A car is accelerated uniformly at 0.15 m per second squared to a speed of 15 m/s.the speed is maintained for 3min,and it is retardaded uniformly to come rest in 8min after starting.draw the velocity time graph and use it to find the average velocity.

1
Expert's answer
2022-05-08T14:29:46-0400

"<v>=\frac s{∆t}=\frac s{\frac{s_1}{v_1}+\frac{s_2}{v_2}}=0 .05~\frac ms."